A beaker with 1.90×102 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 8.40 mL of a 0.410 M HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740.
Answer #1

pH of buffer

pH = pKa + log [salt/[acid]

5.000 = 4.740 + log [salt/[Acid]

Or, log [salt /acid] = 5.000 - 4.740 = 0.26.

Or, [salt/acid] = 100.26 = 1.819.

Given, [ salt ] + [ acid] = 0.100 M

Or, 1.819 [acid] + [acid] = 0.100

Or, 2.819 [acid] = 0.100

Or, [acid] = 0.100/2.819 = 0.0354 M

Then [salt] = (0.100 - 0.0354) = 0.0646 M.

milmoles of acetic acid = 0.0354* 1.90*102 = 6.726

milimoles of salt = 0.0646*1.90*102 = 12.274

milimoles of HCl added = 0.410*8.40 = 3.444

The reaction when HCl is added

CH3COONa + HCl \to CH3COOH + H2O

now ice table is

CH3COONa HCl CH3COOH
Initial milimoles 12.274 3.444 6.726
Change -3.444 -3.444 +3.444
equilibrium milimoles 8.827 0 10.17

pH = pKa + log (8.827/10.17)

Or, pH = 4.74 + log ( 0.8679)

Or, pH = ( 4.74 - 0.0615) = 4.678

So, pH in change by (5.000 - 4.678) = 0.322